# Anisotropies of cosmological gravitational wave backgrounds in non-flat spacetime

### v1

@@ Abstract
Recent reports of stochastic gravitational wave background from four independent pulsar-timing-array collaborations have renewed the interest in the cosmological gravitational wave background (CGWB), which is expected to open a new window into the early Universe. Although the early Universe is supposed to be flat from an inflationary point of view, the cosmic microwave background (CMB) data alone from the Planck satellite measurement prefers an enhanced lensing amplitude that can be explained by a closed Universe. In this paper, we propose an independent method to constrain the early-Universe flatness from the anisotropies of CGWB. Using the generalized harmonic decompositions in the non-flat spacetime, we find CGWBs from different physical mechanisms such as cosmic inflation and phase transitions share the same integrated Sachs-Wolfe (ISW) term but possess different SW terms, which would exhibit different behaviors when including the spatial curvature since the ISW effect is more sensitive to the spatial curvature than the SW effect. Furthermore, we provide the cross-correlations between CGWB and CMB, implying a positive or negative correlation between their SW effect terms depending on the GW mechanisms, which may hint at the sign of $f_{\mathrm{NL}}$ when considering non-Gaussianity contributions to anisotropies.
@@ 1 Introduction
A non-zero spatial curvature $K$ can play a crucial role in the evolution of our Universe, especially the large-scale physics and the topology of the Universe [1]. Recent Planck analysis of cosmic microwave background (CMB) anisotropies has pushed the standard cosmological model to an unprecedented precision [2] with $\Omega_{K}=-0.044^{+0.018}_{-0.015}$ (68%, $Planck$ TT,TE,EE + lowE). Combining the Planck-CMB data with lensing and various BAO measurements can indeed constrain the spatial curvature to an extremely flat value $\Omega_{K}=0.0007\pm 0.0019$ (68%, $Planck$ TT,TE,EE + lowE + lensing + BAO). However, these constraints rely on a specific cosmological model and arrive at an anomalously higher lensing amplitude [3, 4, 5]. Nevertheless, a positive curvature can naturally explain the anomalously higher lensing amplitude and alleviate the tensions with the supernova observations at low redshifts [3, 4, 5, 6] (see [7] for a review of tensions). Nevertheless, a non-flat Universe enhances the discordances with most of the late-time cosmological observables like BAO [4]. This motivates us to develop a new method to constrain the spatial curvature $\Omega_{K}$, and the anisotropies of cosmological gravitational wave backgrounds (CGWB) can be such a candidate.
@@ 3 Angular power spectrum of CGWB and its cross-correlation with CMB
In this Section, we will calculate numerically the angular auto-correlation also known as the angular power spectrum of CGWB, which is sensitive to the initial conditions. In the meantime, we will show the cross-correlation of CGWB with CMB temperature anisotropies.
@@ 4 Conclusion and discussion
Due to the extremely weak interactions of gravitons with ordinary matters, the CGWB serves as a clean probe into the early Universe. Furthermore, the anisotropies in the CGWB can offer additional valuable insights into various physical mechanisms generating the CGWBs, such as alleviating the degeneracy of $F_{\mathrm{NL}}$ in scalar-induced gravitational waves (SIGW) [35]. In this study, we have extended the analysis of CGWB anisotropies to a non-flat background. In addition to the conventional total angular momentum method, a more direct approach is proposed to perform harmonic decompositions on scalar functions and then address the Boltzmann equations in a non-zero spatial curvature background with minimal assumptions. We applied this method to derive the expression of auto-correlation and cross-correlation of GWB and CMB angular power spectra. We have shown that variations in initial conditions can significantly influence the final angular power spectra of different CGWBs, as these conditions can either enhance or suppress the SW effect while keeping the ISW effect the same as each other, which can lead to different behaviors when spatial curvature is non-zero. Therefore it is important to study the case when the initial condition is non-adiabatic [57], which may contribute a very different SW effect.
Additionally, relativistic particles may impact the results [58] due to their effects on scalar perturbations, which can subtly alter the SW and ISW effects. Hence, a different number of relativistic particles can also affect the final angular power spectrum when spatial curvature is non-zero. For the SIGW case, we only consider in this paper the situation calculated in Ref. [34], therefore, some other cases like Ref. [47] need to be considered separately. Furthermore, a recent study [59] on SIGW in non-flat spacetime has indicated a very different shape in the GW energy density spectrum solely due to the curvature effect. As a result, a more systematic analysis of the SIGW case is needed, which will be explored in future studies.
@@ Acknowledgments
This work is supported by the National Key Research and Development Program of China Grants No. 2021YFC2203004, No. 2021YFA0718304, and No. 2020YFC2201501, the National Natural Science Foundation of China Grants No. 12422502, No. 12105344, No. 12235019, No. 12047503, No. 12073088, No. 11821505, No. 11991052, and No. 11947302, the Strategic Priority Research Program of the Chinese Academy of Sciences (CAS) Grant No. XDB23030100, No. XDA15020701, the Key Research Program of the CAS Grant No. XDPB15, the Key Research Program of Frontier Sciences of CAS, the Science Research Grants from the China Manned Space Project with No. CMS-CSST-2021-B01 (supported by China Manned Space Program through its Space Application System),
@@ Appendix A Hierarchy equations
It can also be expressed by performing multipole decomposition separately on $A$ and $B$.

### v2

@@ Abstract
Recent reports of stochastic gravitational wave background from four independent pulsar-timing-array collaborations have renewed the interest in the cosmological gravitational wave background (CGWB), which is expected to open a new window into the early Universe. Although the early Universe is supposed to be extremely flat from an inflationary point of view, the cosmic microwave background (CMB) data alone from the Planck satellite measurement prefers an enhanced lensing amplitude that can be explained by a closed Universe. In this paper, we propose an independent method to constrain the early-universe flatness from the anisotropies of CGWB. Using the generalized harmonic decompositions in the non-flat spacetime, we find CGWBs from different physical mechanisms such as cosmic inflation and phase transitions share the same integrated Sachs-Wolfe (ISW) term but possess different SW terms, which would exhibit different behaviors when including the spatial curvature since the ISW effect is more sensitive to the spatial curvature than the SW effect. Furthermore, we provide the cross-correlations between CGWB and CMB, implying a positive or negative correlation between their SW effect terms depending on the GW mechanisms, which may hint at the sign of $f_{\mathrm{NL}}$ when considering non-Gaussianity contributions to anisotropies.
@@ 1 Introduction
A non-zero spatial curvature $K$ can play a crucial role in the evolution of our Universe, especially the large-scale physics and the topology of the Universe [1]. Recent Planck analysis of cosmic microwave background (CMB) anisotropies has pushed the standard cosmological model to an unprecedented precision [2] with $\Omega_{K}=-0.044^{+0.018}_{-0.015}$ (68%, $Planck$ TT,TE,EE + lowE). Combining the Planck-CMB data with lensing and various BAO measurements can indeed constrain the spatial curvature to an extremely flat value $\Omega_{K}=0.0007\pm 0.0019$ (68%, $Planck$ TT,TE,EE + lowE + lensing + BAO). However, these constraints rely on a specific cosmological model and arrive at an anomalously higher lensing amplitude [3, 4, 5]. Nevertheless, a positive curvature can naturally explain the anomalously higher lensing amplitude and alleviate the tensions with the supernova observations at low redshifts [3, 4, 5, 6] (see [7] for a review of tensions). Nevertheless, a non-flat Universe enhances the discordances with most of the late-time cosmological observables like BAO [4]. In addition to this, one cannot simply assume $\Omega_{K}=0$ [8]. These motivate us to develop a new method to constrain the spatial curvature $\Omega_{K}$, and the anisotropies of cosmological gravitational wave backgrounds (CGWB) can be such a candidate.
@@ 3 Angular power spectrum of CGWB and its cross-correlation with CMB
In this Section, we will calculate numerically the angular auto-correlation also known as the angular power spectrum of CGWB, which is sensitive to the initial conditions. In the meantime, we will show the cross-correlation of CGWB with CMB temperature anisotropies. At last, we will briefly analyze the signal-to-noise ratio (SNR) for anisotropies from FOPTs.
@@ 3.3 Cross-correlations between CGWB and CMB
3.4 SNR of anisotropies from FOPTs
To relate to the experiments, an analysis of SNR is needed. We will follow the definition of Refs. [21, 48, 58], and the SNR is related to the multipole $\mathrm{SNR}_{l}^{2}$ via $\mathrm{SNR}^{2}=\sum_{l}(2l+1)\mathrm{SNR}_{l}^{2}$, where the multipole SNR, $\mathrm{SNR}_{l}$, is defined as
$\mathrm{SNR}_{l}^{2}=\int\mathbf{d}fC_{l}^{\mathrm{GW}}(f)\left(\frac{\Omega_{\mathrm{GW}}(f)}{\Omega_{\mathrm{GW},n}^{l}(f)}\right)^{2}.$ (3.26)
Here, $\Omega_{\mathrm{GW},n}^{l}(f)$ is the effective angular sensitivity of the detector network to the $l$-th multipole,
$\Omega_{\mathrm{GW},n}^{l}(f)^{-2}\equiv T_{\mathrm{obs}}\sum_{AB}\left(\frac{2}{5}\right)^{2}\left(\frac{4\pi^{2}f^{3}}{3H_{0}^{2}}\right)^{-2}\frac{\sum_{m}|\mathcal{A}_{AB}^{lm}{f}|^{2}}{N_{f}^{A}N_{f}^{B}(2l+1)},$ (3.27)
where $T_{\mathrm{obs}}$ is the time of observation, $N_{f}^{A}$ is the noise power spectrum density for detector A, and $\mathcal{A}_{AB}^{lm}$ denotes the spherical harmonic transform of the antenna pattern for the detector pair $AB$. In the following, we will use the code schNell [58] to calculate $\Omega_{\mathrm{GW},n}^{l}(f)$. As for the numerical part, we choose the observation time $T_{\mathrm{obs}}$ to be 4 years and fix the pivot frequency to be 1 mHz in the code. We adopt the effective sensitivity of the detector network LISA  [^2] This result is marginally consistent with the analysis of Ref. [21], where the discontinuity of $l=1$ curve around 3 mHz might be numerical instability but it is unharmful to our calculations after integrations. as shown in Fig. 4, and one can see the sensitivity declines quickly when $l$ gets larger.
Figure 4: The effective sensitivity of the detector network LISA, for multipoles up to $l=10$. Even (odd) multipoles are shown with solid (dashed) lines.
Recalling that although our analysis above this subsection is general, we have specifically adopted an illutrative value for $n_{\mathrm{GW}}$ (3.16) from the recent fitting result of PTA data [56], and hence all numerical results above this subsection are subjected to the frequency band of $\mathcal{O}(\mathrm{nHz})$  [^3] For those who are interested in the PTA sensitivity of anisotropies, please refer to Refs. [59, 60], and for those who are interested in other detectors or frequency bands, please refer to Ref. [61].. Nevertheless, our analysis can easily be extended to the frequency band of $\mathcal{O}(\mathrm{mHz})$ given a specific model of GW energy spectrum at that frequency band. In this subsection, we will roughly estimate the SNR for the case with a FOPT using the LISA detector network, while generalizing to other cases is straightforward. As for the phase-transition case, we consider the total GW spectrum $\Omega_{\mathrm{GW}}$ is dominated by the sound-wave spectrum $\Omega_{\mathrm{sw}}$ fitted by the LISA Cosmology Working Group (see, for example, Ref. [62]),
$\Omega_{\mathrm{GW}}^{\mathrm{PT}}(f)h^{2}=AS_{\mathrm{sw}}(f/f_{p}),$ (3.28)
where $f_{p}$ refers to the peak frequency, and the shape function is fitted as
$S_{\mathrm{sw}}(f/f_{p})\equiv\left(\frac{f}{f_{p}}\right)^{3}\left(\frac{7}{4+3(f/f_{p})^{2}}\right)^{7/2}.$ (3.29)
We will choose $A=10^{-12}$ to be consistent with numerical results from Refs. [63, 53], and the peak frequency is fixed at $f_{p}=3\times 10^{-3}$z, which is shown in the left panel of Fig. 5. In this case, $n_{\mathrm{GW}}$ is a frequency-dependent quantity as shown in the right panel of Fig. 5, which asymptotes to constant powers at lower and higher frequencies.
Before we calculate $\mathrm{SNR}_{l}$, we can visualize the relative magnitudes of $\Omega_{\mathrm{GW}}^{\mathrm{PT}}$ with respect to the most optimistic sensitivity $\Omega_{\mathrm{GW}}^{l=0}$ as shown in Fig. 6, where we have multiplied the monopole part of $\Omega_{\mathrm{GW}}$ with $\sqrt{C_{l}}$ as appear in the integrand of definition of SNR (3.26). The numerical result of each $\mathrm{SNR}_{l}$ is presented in the left panel of Fig. 7 for different choices of spatial curvature, which is far below the current detection ability of LISA. To obtain $\mathcal{O}(1)$ SNR, we can raise the amplitude $A$ of $\bar{\Omega}_{\mathrm{GW}}$ up to $\mathcal{O}(10^{-6})$. Therefore, it will be quite challenging for LISA to detect the corresponding SNR, which motivates us to look at the cross-correlation, whose SNR is defined as [57]
$(\mathrm{SNR}^{c})^{2}=f_{\mathrm{sky}}\sum_{l=2}^{l_{\mathrm{max}}}(2l+1)\frac{(C_{l}^{\mathrm{CMB}\times\mathrm{GWB}})^{2}}{(C_{l}^{\mathrm{CMB}\times\mathrm{GWB}})^{2}+(C_{l}^{\mathrm{GWB}}+N_{l}^{\mathrm{GWB}})C_{l}^{\mathrm{CMB}}}\equiv f_{\mathrm{sky}}\sum_{l=2}^{l_{\mathrm{max}}}(2l+1)(\mathrm{SNR}_{l}^{c})^{2}.$ (3.30)
Here, $f_{\mathrm{sky}}$ denotes the sky coverage that can be simply set to $1$ for a full-sky coverage, and we also neglect the noise of CMB instruments since it is much smaller than $C_{l}^{\mathrm{CMB}}$ at low $l$. The $l$-th multipole noise power spectrum $N_{l}^{\mathrm{GWB}}$ can be computed as  [^4] Note the cosmic variance is not included for both $N_{l}^{\mathrm{GWB}}$ and $N_{l}^{\mathrm{CMB}}$ for an order-of-magnitude estimation.
$N_{l}^{\mathrm{GWB}}=\left[\int\mathrm{d}f\frac{\bar{\Omega}_{\mathrm{GW}}(f)^{2}}{\Omega_{\mathrm{GW,n}}^{l}(f)^{2}}\right]^{-1}.$ (3.31)
In the right panel of Fig. 7, we have shown the numerical results of $\mathrm{SNR}^{c}$, which is much larger than the $\mathrm{SNR}$ of auto-correlation. Again, if the amplitude $A$ of the GW power spectrum can reach up to $\mathcal{O}(10^{-9})$, the SNR of cross-correlation would arrive at $\mathcal{O}(1)$, which is consistent with Ref. [57].
Figure 5: Left: The gravitational wave energy density spectrum of phase transition, and the sensitive curve of LISA detector is also given. The amplitude $A$ is chosen as $A=10^{-12}$, which is consistent with the numerical result of [63, 53], and the peak frequency is set to be $f_{p}=3\times 10^{-3}$z. Right: The tensor spectrum tilde $n_{\mathrm{Gw}}$ is given, and one can see it depends on frequency in this situation. If the peak frequency shift to other frequency band, one will get a constant $n_{\mathrm{Gw}}$ in this $\mathcal{O}$(mHz) frequency band.
Figure 6: The strength of PT signal $\sqrt{C}_{l}\bar{\Omega}_{\mathrm{GW}}h^{2}$ and the sensitivity curve of LISA detector.
Figure 7: Left: The $\mathrm{SNR}_{l}$ of the phase transition mechanism. Right: The $\mathrm{SNR}^{c}$ of the cross-correlation in the phase transition case.
@@ 4 Conclusion and discussion
Due to the extremely weak interactions of gravitons with ordinary matters, the CGWB serves as a clean probe into the early Universe. Furthermore, the anisotropies in the CGWB can offer additional valuable insights into various physical mechanisms generating the CGWBs, such as alleviating the degeneracy of $F_{\mathrm{NL}}$ in scalar-induced gravitational waves (SIGW) [36]. In this study, we have extended the analysis of CGWB anisotropies to a non-flat background. In addition to the conventional total angular momentum method, a more direct approach is proposed to perform harmonic decompositions on scalar functions and then address the Boltzmann equations in a non-zero spatial curvature background with minimal assumptions. We applied this method to derive the expression of auto-correlation and cross-correlation of GWB and CMB angular power spectra. We have shown that variations in initial conditions can significantly influence the final angular power spectra of different CGWBs, as these conditions can either enhance or suppress the SW effect while keeping the ISW effect the same as each other, which can lead to different behaviors when spatial curvature is non-zero. Therefore it is important to study the case when the initial condition is non-adiabatic [64] or isocurvature [65], which may contribute a very different SW effect. We have also prospected the detection ability of LISA for the GW anisotropies from a FOPT.
Additionally, relativistic particles may also impact the results [66] due to their effects on scalar perturbations when the anisotropic perturbation is non-zero (namely $\Phi\neq\Psi$), which can subtly alter the SW and ISW effects, otherwise, the result will not change [65]. Hence, a different number of relativistic particles can also affect the final angular power spectrum when the spatial curvature is non-zero. For the SIGW case, we only consider in this paper the situation calculated in Ref. [35], therefore, some other cases like Ref. [48] need to be considered separately. Furthermore, a recent study [67] on SIGW in non-flat spacetime has indicated a very different shape in the GW energy density spectrum solely due to the curvature effect. As a result, a more systematic analysis of the SIGW case is needed, which will be explored in future studies.
@@ Acknowledgments
This work is supported by the National Key Research and Development Program of China Grants No. 2021YFC2203004, No. 2021YFA0718304, and No. 2020YFC2201501, the National Natural Science Foundation of China Grants No. 12422502, No. 12105344, No. 12235019, No. 12047503, No. 12073088, No. 11821505, No. 11991052, and No. 11947302, the Strategic Priority Research Program of the Chinese Academy of Sciences (CAS) Grant No. XDB23030100, No. XDA15020701, the Key Research Program of the CAS Grant No. XDPB15, the Key Research Program of Frontier Sciences of CAS, the Science Research Grants from the China Manned Space Project with No. CMS-CSST-2021-B01 (supported by China Manned Space Program through its Space Application System). Our code is available at this Github repository.
@@ Appendix A Hierarchy equations
It can also be expressed by performing multipole decomposition separately on $A$ and $B$,